Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

pm1weekly_2004-101_2004-107

Mission Operations Days: 2004/101 to 2004/107
April 10, 2004 00:00:00 GMT to April 16, 2004 00:00:00 GMT


Aqua Spacecraft and MODIS Instrument Status:

PM-1 "Aqua" is in Fine Pointing Normal Mode
FM1 MODIS is in B-Side Nominal Science Mode
ComponentState(s)Comment
BlackbodyA Off; B On (285K)Nominal Science Mode
Calibration ElectronicsA Off; B OnNominal Science Mode
Control ProcessorA Off; B OnNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ClosedNominal Science Mode
FDDI FormatterA Off; B OnNominal Science Mode
FDDI PortA Off; B OnNominal Science Mode
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 Off, #2 OnNominal Science Mode
PV VIS FPAA Off; B OnNominal Science Mode
PV NIR FPAA Off; B OnNominal Science Mode
PV SMIR FPAA Off; B OnNominal Science Mode
PV LWIR FPAA Off; B OnNominal Science Mode
PC LWIR FPAA Off; B OnNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOffNominal Science Mode
RC SMIR FPA HeaterOn (83K)Nominal Science Mode
Scan AssemblyA Off; B OnNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeatersEnabledNominal Science Mode
PS2 Survival HeatersEnabledNominal Science Mode
Timing GeneratorA Off; B OnNominal Science Mode
Flight SoftwareRev BJ + 1 set of patches
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
Nominal Science Mode
Inhibit Ids SetNoneNominal Science Mode

This Week’s Completed MODIS Activities:

Saturday, April 10, 2004

None

Sunday, April 11, 2004

None

Monday, April 12, 2004

2004/103 21:04:53-21:18:51 MCL MOD_OA16_SDSM_SCREEN.2 OA-16: SD/SDSM Screened - B-Side
2004/103 22:43:47-22:57:45 MCL MOD_OA15_SDSM_OPEN.1 OA-15: SD/SDSM Open - B-Side
2004/103 23:40:00-23:40:16 MCL MOD_SRCA_SETUP_NO_L2.2 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/103 23:43:26-00:17:38 MCL MOD_OA19_SRCA_FULL_RAD.1 OA-19: SRCA Full Radiometric Calibration - B Side
Data Collect (23:59:20-00:17:38) Centered on S/C Night (23:51:34-00:25:23)

Tuesday, April 13, 2004

None

Wednesday, April 14, 2004

None

Thursday, April 15, 2004

None

Friday, April 16, 2004

None

This Week’s Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Calibrations for April will include biweekly SDSM, monthly SRCA Full Radiometric, bimonthly PV Ecal, and quarterly SRCA Full Spectral and Black Body calibrations.

Friday, April 23, 2004

2004/114 02:00:00 MCL MOD_BB_TEMP_SET_270K_B.1
BB Temp Set to 270K for CP B Coefficients
2004/114 20:00:00 MCL MOD_BB_TEMP_SET_280K_B.1
BB Temp Set to 280K for CP B Coefficients

Maneuvers:

The next Aqua MODIS Lunar Calibration Roll will be scheduled for 2004/121 (April 30, 2004). Data collect time is 17:12:38 to 17:19:08 GMT. The roll angle will be -6.9795 degrees.

No official date has yet been scheduled for the Aqua Deep Space Calibration pitch maneuver(s).

MODIS Anomalies:

None

General Instrument Comments:

FM1 MODIS is in Science Mode. The SVD and NAD are Open, the survival heaters are enabled, and the Black Body is heated to 285 K.

MODIS Telemetry Trends:

The FM1 MODIS instrument telemetry is trending nominally.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links webpages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 229.8 of 500 hours
SRCA 10W Lamp #2: 184.3 of 500 hours
SRCA 10W Lamp #3: 195.8 of 500 hours
SRCA 10W Lamp #4: 69.0 of 500 hours

SRCA 1W Lamp #1: 507.3 of 5000 hours
SRCA 1W Lamp #2: 269.8 of 5000 hours

Solar Diffuser Door: 2194 of 3022 Movements
Nadir Aperture Door: 1053 of 1316 Movements
Space View Door: 632 of 1316 Movements